OpenStudy (anonymous):

Find all the solutions of the equation 3sin^2(x)-7sin(x)+2=0. The answers is x1=_______, x2=__________. x1

OpenStudy (campbell_st):

this is basically a quadratic equation that can be factored any ideas..?

OpenStudy (anonymous):

(3sin(x)-1)(sin(x)-2)=0. Is that right?

OpenStudy (campbell_st):

thats correct... so you need to solve 3sin(x) - 1 = 0 and sin(x) - 2 = 0 so what solutions do you get....in the form sin(x) =

OpenStudy (anonymous):

sin(x)=1/3 and 2

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

great... so what do you know about sin(x) = 2

OpenStudy (anonymous):

It's no solution right?

OpenStudy (campbell_st):

that's correct, it's undefined so you only case is sin(x) = 1/3 so there are solutions in the 1st and 2nd quadrants

OpenStudy (anonymous):

I don't know how to find those solutions. @campbell_st

OpenStudy (campbell_st):

well I'd probably express it ad decimal radians find \[x = \sin^{-1}(\frac{1}{3})\] so 1st quadrant angle is x 2nd quadrant angle is 180 - x